OpenStudy (anonymous):

4. Given the linear equation y = x – 7, find the y-coordinates of the points (-2, ), (7, ), and (2, ). Please show all of your work. Plot these points and graph the linear equation

OpenStudy (anonymous):

replace x by -2 and see what you get for y. etc. so the first one is \[y=-2-7=-9\] and the coordinate is therefore \[(-2,-9)\]

OpenStudy (anonymous):

for x = 7 write \[y=7-7=0\] so you have \[(7,0)\] can you do the last one?

OpenStudy (anonymous):

how do you plot them on a graph

OpenStudy (anonymous):

how did you get the 7-7

OpenStudy (anonymous):

y=2-7=-5 =2,-5